OpenStudy (anonymous):

z varies directly with x and inversely with y. When x=6 and y=2, z=15. Find z when x=4 and y=9

OpenStudy (anonymous):

Can someone explain how to do this?

OpenStudy (mertsj):

start out with z=

OpenStudy (mertsj):

Then put the directly stuff on top and the inversely stuff on the bottom.

OpenStudy (anonymous):

I made a proportion of 6/15 4/z

OpenStudy (anonymous):

I tried to cross multiply and divide and got 10

OpenStudy (mertsj):

Then multiply by k for the constant.

OpenStudy (anonymous):

How do you determine the constant?

OpenStudy (mertsj):

\[z=\frac{kx}{y}\] \[\frac{zy}{x}=k\] \[\frac{15(2)}{6}=\frac{9z}{4}\]

OpenStudy (mertsj):

Here is how to find k Replace z with 15, y with 2 and x with 6 and solve for k

OpenStudy (anonymous):

2.2 repeating?

OpenStudy (anonymous):

Is that what k equals?

OpenStudy (mertsj):

\[15=\frac{6k}{2}\] \[30=6k\] \[5=k\]

OpenStudy (anonymous):

Isn't it 15 times 2?

OpenStudy (mertsj):

Yes. and 15 times 2 = 30

OpenStudy (anonymous):

and isn't it 9k/4

OpenStudy (mertsj):